# Exploring nonlinearities in a positive ion-negative ion (PINI) plasma: can other processes mimic debris-induced effects?

### v3

@@ Abstract
In this work, an analysis of nonlinear waves and structures induced by an external charged debris in a positive ion-negative ion (PINI) plasma is presented. The results obtained are compared with findings from available experiments involving PINI plasma. The process of formation of different nonlinear structures is examined theoretically through a forced Koretweg-de Vries (fKdV) equation, which is also verified with a multi-fluid flux-corrected transport simulation code (mFCT). Various processes which are responsible for different nonlinear waves and structures excited by differently charged external debris are pointed out. This work also points out the similarities in different nonlinear structures excited by an external charged debris and the underlying processes (this work) and those observed experimentally through processes that do not involve any external debris.
@@ I Introduction
Nonlinear waves ubiquitously drive the dynamics of nature’s most intricate systems, shaping complex phenomena from plasma instabilities to oceanic rogue waves in ways both beautiful and unpredictable. Plasma solitons and shocks remain prime examples of such structures that govern the energy transport and wave dynamics both in laboratory and astrophysical environments. Solitons or solitary waves can be considered as localized compressions or rarefactions of plasma densities arising out of the interplay between nonlinearity and dispersion. On the contrary, shock waves are usually associated with some dissipation mechanisms such as kinematic viscosity, Landau damping, and collision among ions and neutrals along with inherent nonlinearity, dispersion, and weak dissipation. During the nonlinear evolution of ion-acoustic wave (IAW) in a dissipative plasma, the leading edge of a propagating wave gets steepened as dissipation dominates over dispersion and a shock front is formed. In contrast to this, in a collision-less dispersive fluid, shock waves can form without dissipation and often have an oscillatory signature and are known as dispersive shock waves (DSWs) [1, 2, 3, 4]. Nevertheless, the nature of these nonlinear structures is largely determined by the background plasma conditions, constituents, and the associated physical parameters.
@@
The characteristics of ion-acoustic solitons or shock waves in a usual electron-ion ($e$-$i$) plasma are significantly different from that of a multi-ion plasma such as a positive ion-negative ion (PINI) plasma [5, 6, 7, 8]. In fact, an $e$-$i$ plasma can only support compressive solitons with positive electrostatic potential. However, with an addition of even a small amount of one or more extra ion species to the plasma, drastically changes the dynamics of the plasma. It is observed that in a multicomponent plasma, the IAW propagation is mainly influenced by the lighter ions and changes in their concentrations and temperatures may lead to substantial modification of the wave structure via the plasma potential [9, 10, 11]. If the added ion species is negatively ionized, then after a critical concentration of these negative ions, the plasma potential can even change its polarity to form a negative or rarefactive soliton [10, 11]. When an ion beam is injected into such a plasma, beam-plasma interaction makes the evolution even more complex in terms of nonlinear wave structures compared to a simple plasma [11, 12, 13]. For this reason, investigations on the plasmas with negative ions as one of the constituents has always been a topic of interest. Plasmas in astrophysical and space environments such as cometary tails, planetary ionospheres, and interstellar and molecular clouds usually are inherently multi-species with two or more than two types of ions and offer a rich platform for studying a variety of wave motions with subtle intricacies [14, 15]. These multi-species plasmas with negative ions are also easily produced in a laboratory and are well studied owing to their vital role in major scientific and technological applications such as plasma processing or fusion plasmas [8]. However, the response of a PINI plasma to an external charged debris is a topic yet to be explored fully. Moreover, as we explain in the next subsection, we raise a fundamental question about plasma-debris interaction – can other nonlinear processes mimic debris-induced nonlinearities? As we shall see, under certain circumstances, this is indeed true. So, we believe that an investigation of nonlinear wave propagation in a PINI plasma embedded with an external charged debris is definitely a topic worth studying, which should help us in understandings about the intricacies of the processes leading to formation of these nonlinear structures in multicomponent plasma.
@@ II Plasma model and governing equations
The normalized electron density is given by $n_{e}=e^{\phi}$.
@@ II.1 The morphology of debris-induced nonlinearities
Before we investigate the nonlinear waves and structures induced by a charged debris in a PINI plasma, it is of relevance to discuss a bit about the fundamental physics of these nonlinearities in an $e$-$i$ plasma. We now know that the response of an $e$-$i$ plasma to an embedded charged debris very well depends on the nature of the charge of the debris [22]. It has been only recently shown that the formation of bright pinned solitons in the ion-acoustic regime due to the presence of negatively charged debris is basically a manifestation of the trapped ions in phase-space vortices formed due to ion-ion counter streaming instability (IICSI) [19]. Besides, proper spatial resolution of the pinned solitons requires a certain relative velocity between the debris and the plasma, though higher relative velocities may cause the pinned solitons to disappear [17, 19]. In contrast, a positively charged debris creates an ion hole, exciting an ion-acoustic wave (IAW), which propagates away from the site of the debris. However, when there is a relative motion between the debris and the plasma, the IAW becomes a dispersive shock wave (DSW) in the precursor region [22].
@@ II.2 Mathematical reduction
We note that the above plasma model can be analyzed both mathematically and numerically. Typically, in these class of problems, one uses a scale transformation and reduces it to a pseudo potential (or Sagdeev potential) form [25] through the Poisson equation, which guarantees a soliton solution, though not necessarily a ‘${\rm sech}$’-type solution as one would have obtained through a KdV equation [26]. The other two approaches would be a reductive perturbation analysis to reduce the system to a KdV equation or a nonlinear Schrödinger equation (NLSE) [27]. However, with an external charged debris term, the model is simply not amenable to pseudo-potential analysis and that leaves us with other two options.
Employing the reductive perturbation method with the external charged debris term, we end up with either a forced-KdV (fKdV) or a forced-NLSE equation. At this point, we should emphasize that numerically we are going to a use flux-corrected transport (FCT) method (as mentioned before) to simulate the dynamics, which is very efficient in detecting sharp discontinuities that can result from the debris-induced perturbation. Also the FCT formalism does not employ any approximations and solves the full set of equations, which in principle should provide us with both KdV-type as well as NLSE-type modulated wave solutions. However, the single hump perturbation that we expect from an external charged debris, the numerical solution is most likely to favor KdV-type solution which is also closer to the real-world and observable dynamics. In what follows, we shall therefore analyze the above model mathematically, using the fKdV equation only.
@@ III Forced-KdV dynamics
where $\varepsilon$ is the expansion parameter, which is $\ll 1$. One question related to the order of perturbation for the external charged debris may naturally arise here. In this context, two points to be noted here – that we use the KdV equation for weak nonlinearity only and how strong a perturbation an external charged debris may induce in the bulk plasma. Our assumption of using a second order perturbation for the debris is quite in line with the weak nonlinearity assumption on the basis of which the KdV equation is derived. As far as the strength of perturbation is concerned, we shall see in subsequent sections that the experimentally observed nonlinearities are quite close to our numerical solutions (obtained via FCT simulation) as well as the fKdV solutions, which demonstrates that our assumption of introducing the external charged debris as second order perturbation is indeed consistent. Besides, using a second order perturbation for debris leaves us with the simplest form of fKdV equation.
@@
As mentioned before, we now provide an explanation about our ‘dissipation-less’ assumption of the plasma model. The dissipation in this case may come in two flavors – through collisions and Landau damping. There are examples of work involving multicomponent plasma (not necessarily PINI plasma) with collisions as well as with Landau damping term [28]. While the inclusion of different collisional terms in the fluid model is quite straight forward [29, 30], the same is not true in the case of Landau damping. However, as shown by VanDam [31], it is indeed possible to include a dissipation-like term arising out of wave-particle resonance (Landau damping) in a KdV-like formalism. In Section III-A.1, while comparing our fKdV solutions with ion-beam-induced nonlinearities, we shall see that the $e$-$i$ collision time $\tau_{ei}\sim 3.5\,{\rm milli}\,{\rm second}$. However, the ion-acoustic time scale, which is the characteristic time scale of this work, is $\tau_{{\rm IA}}\sim 0.03\,{\rm milli}\,{\rm second}$. So, we can probably safely neglect collisional dissipation. Coming back to now Landau damping, VanDam [31] has shown that a dissipation-like term can be incorporated into the KdV equation with the linear Landau damping rate $\gamma=\tilde{\gamma}|k|$ (see Appendix A) where
@@ III.1 Positively charged external debris
where $e$ is the electronic charge and $R$ is the radius of a spherical debris [32]. This result is rigorously proved for impurity particles in a plasma using the orbit motion limited (OML) theory [33] and numerical simulation, which can be readily interpreted for charged debris as well. This can also be extended to debris with positive charge.
Figure 1: Nonlinear wave for a flowing PINI plasma with a positively charged external debris. All panels are drawn in the rest frame of the debris, which is in the middle of a panel, denoted by a vertical dashed line. The shaded vertical regions indicate the extent of the charged debris. The shaded envelopes on the right hand panels indicated the envelopes of the corresponding dark pinned solitons.
If we now consider the case of charged debris in space plasma in low earth orbit (LEO), the size of the debris can vary from being very small to very large, but anything larger than a few Debye length is capable of inducing nonlinear structures. Considering typical parameters for LEO plasma [34, 35], the value for $\beta_{e}$ comes out to be $\sim 10^{-6}$ even for $R\sim\lambda_{D}$. For ions (positively charged debris), this comes out to be $\beta_{i}\sim 10^{-3}$. In laboratory experiments involving debris $\beta_{e}\sim 10^{-5}$ assuming debris of size $\sim\lambda_{D}$ [36].
Figure 2: Nonlinear wave for a flowing plasma with a negatively charged external debris. All other particulars are same as in Fig.1. The shaded vertical regions indicate the extent of the negatively charged debris.
These results are entirely in conformity with what we know about formation of debris-induced pinned solitons and DSWs. We now know that in an $e$-$i$ plasma the pinned solitons formed due to moving negatively charged debris is basically a manifestation of the ions trapped in the phase space vortices formed by the IICSI [19], whereas a positively charged debris induces a DSW in the precursor region. So, for $\delta\sim 1$ (vey less concentration of negative ions), the plasma behaves like the usual $e$-$i$ plasma and no pinned solitons can form for a moving positively charged debris.
However for $\delta>1$, the negative ions are attracted by the positively charged debris causing a IICSI-like phenomena for negative ions, leading to the formation of phase space vortices which in turn cause formation of dark pinned solitons at the debris position.
@@ III.1.1 Comparison with ion-beam-induced nonlinearities
Let us now discuss the first case where under certain circumstances, a positive ion-beam in a PINI plasma mimics the responses of a debris-induced nonlinearities. In this experiment, the positive ion-beam is produced in situ by accelerating a part of the ion population in the plasma so that overall quasi-neutrality is maintained throughout [11]. This case is also interesting as there are several investigations being carried out by different authors to analyze the formation of nonlinear structures in such plasmas [12, 13].
Figure 3: FCT simulation results of propagation of nonlinear wave in a PINI plasma with and embedded external positively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.1. The left panel (a) in the above figure corresponds to Fig.1(c) and the right panel (b) above corresponds to Fig.1(d). These results are plotted at the and of $\tau=100\equiv\varepsilon^{3/2}t$.
Let us consider such a three-component PINI plasma, consisting of positive ion $(n_{+})$, negative ion $(n_{-})$, and electron $(n_{e})$, in the presence of a positive ion beam $(b^{+})$, as considered in the experimental work by Nakamura [11] (N99 hereafter). Note that the whole system is to be considered quasi-neutral,
@@
where $\ln\Lambda$ is the Coulomb logarithm $\sim 10-12$. For $n_{e}\sim 10^{7}\,{\rm cm}^{-3}$ [11], we have $\tau_{ei}\sim 3.5$ milli second. We now argue that as $\tau_{ei}\gg\tau_{{\rm IA}}$, it can be safely assumed that in the ion-acoustic timescale, the plasma will not be able to thermalize. As a result, a positive ion beam propagating through the plasma will not be able to thermalize with the bulk plasma and for all practical purposes, the beam will behave as an external debris. It has been observed in N99 that below a certain critical value for beam velocity, increasing $r$ (higher $\delta$) causes the nonlinearity to become negative from positive. However, beyond the critical value of the beam velocity, the nonlinearity remains positive irrespective of the initial pulse (which is reported only at higher $\delta$ in N99). In our case with debris as well, we see a similar behavior as evident from the panels in the left column of the figure i.e. Fig.1(a), (b), and (c). We can see that when debris velocity is $\apprle 1.1$, for low $\delta$ $(\sim 1)$, there is a steepening in the leading edge indicating positive nonlinearity while for higher $\delta$, the steepening is on the falling edge, which indicates negative nonlinearity. We note that the ‘positive nonlinearity’ in the case of a compressive or bright soliton is referred with respect to the nonlinear term $A\phi^{(1)}\partial_{\xi}\phi^{(1)}$ [see Eq.(14)] in a usual KdV equation (without the debris term), which is responsible for steepening of the leading edge of a soliton. On the other hand, when the steepening occurs in the falling edge, it is usually termed as ‘negative nonlinearity’. When debris velocity increases [Fig.1(f)], the negative nonlinearity which was observed for say $\delta=1.5$ and with low $v_{d}$, decreases and get balanced by dispersion giving rise to a bright KdV soliton. These results are also confirmed through our FCT simulation, which are being presented in Section IV.

### v4

@@ Abstract
In this work, an analysis of nonlinear waves and structures induced by an external charged debris in a positive ion-negative ion (PINI) plasma is presented. The results obtained are compared with findings from available experiments involving PINI plasma. The process of formation of different nonlinear structures is examined theoretically through a forced Korteweg-de Vries (fKdV) equation, which is also verified with a multi-fluid flux-corrected transport simulation code (mFCT). Various processes which are responsible for different nonlinear waves and structures excited by differently charged external debris are pointed out. This work also points out the similarities in different nonlinear structures excited by an external charged debris and the underlying processes (this work) and those observed experimentally through processes that do not involve any external debris.
@@ I Introduction
Nonlinear waves ubiquitously drive the dynamics of nature’s most intricate systems, shaping complex phenomena from plasma instabilities to oceanic rogue waves in ways both beautiful and unpredictable. Plasma solitons and shocks remain prime examples of such structures that govern the energy transport and wave dynamics both in laboratory and astrophysical environments. Solitons and solitary waves can be considered as localized compressions or rarefactions of plasma densities arising out of the interplay between nonlinearity and dispersion. On the contrary, shock waves are usually associated with some dissipation mechanisms such as kinematic viscosity, Landau damping, and collision among ions and neutrals along with inherent nonlinearity, dispersion, and weak dissipation. During the nonlinear evolution of ion-acoustic wave (IAW) in a dissipative plasma, the leading edge of a propagating wave gets steepened as dissipation dominates over dispersion and a shock front is formed. In contrast to this, in a collision-less dispersive fluid, shock waves can form without dissipation and often have an oscillatory signature and are known as dispersive shock waves (DSWs) [1, 2, 3, 4]. Nevertheless, the nature of these nonlinear structures is largely determined by the background plasma conditions, constituents, and the associated physical parameters.
@@
The characteristics of ion-acoustic solitons or shock waves in a usual electron-ion ($e$-$i$) plasma are significantly different from that of a multi-ion plasma such as a positive ion-negative ion (PINI) plasma [5, 6, 7, 8]. In fact, an $e$-$i$ plasma can only support compressive solitons with positive electrostatic potential. However, with an addition of even a small amount of one or more extra ion species to the plasma, drastically changes the dynamics of the plasma. It is observed that in a multicomponent plasma, the IAW propagation is mainly influenced by the lighter ions and changes in their concentrations and temperatures may lead to substantial modification of the wave structure via the plasma potential [9, 10, 11]. If the added ion species is negatively ionized, then after a critical concentration of these negative ions, the plasma potential can even change its polarity to form a negative or rarefactive soliton [10, 11]. When an ion beam is injected into such a plasma, beam-plasma interaction makes the evolution even more complex in terms of nonlinear wave structures compared to a simple plasma [11, 12, 13]. For this reason, investigations on the plasmas with negative ions as one of the constituents have always been a topic of interest. Plasmas in astrophysical and space environments such as cometary tails, planetary ionospheres, and interstellar and molecular clouds usually are inherently multi-species with two or more than two types of ions and offer a rich platform for studying a variety of wave motions with subtle intricacies [14, 15]. These multi-species plasmas with negative ions are also easily produced in a laboratory and are well studied owing to their vital role in major scientific and technological applications such as plasma processing or fusion plasmas [8]. However, the response of a PINI plasma to an external charged debris is a topic yet to be explored fully. Moreover, as we explain in the next subsection, we raise a fundamental question about plasma-debris interaction – can other nonlinear processes mimic debris-induced nonlinearities? As we shall see, under certain circumstances, this is indeed true. So, we believe that an investigation of nonlinear wave propagation in a PINI plasma embedded with an external charged debris is definitely a topic worth studying, which should help us in understandings about the intricacies of the processes leading to formation of these nonlinear structures in multicomponent plasma.
@@ II Plasma model and governing equations
The normalized electron density is given by $n_{e}=e^{\phi}$. It should however be noted that with an external charged debris, the above quasi-neutrality condition does not hold good at the site of debris and this is also the precise reason why the system becomes non-integrable (see the next Section for more discussion). We also note that given the case for ion-acoustic solitons or solitary waves, instantaneously there may be a violation of the Boltzmannian electrons when $e\phi\sim m_{e}v_{{\rm th}e}$, where $v_{{\rm th}e}$ is the electron thermal velocity. However, considering the overall IA timescale $\tau_{\rm IA}\gg\omega_{pe}^{-1}$, the electron plasma frequency, we can justify the electron equilibrium to be quasi-static even though the ion response is nonlinear. This can even be justified from kinetic theory as well by using the stationary Vlasov equation for electrons (dimensional)
$v_{e}\frac{\partial f_{e}}{\partial x}-\frac{e}{m_{e}}\frac{\partial\phi}{\partial x}\frac{\partial f_{e}}{\partial v_{e}}=0,$ (6)
assuming $\partial f_{e}/\partial x\sim 0$ in the IA timescale, which leads to the Boltzmannian distribution on integration [25]
[…]
In the above expressions, $v_{e},m_{e}$ are respectively electron velocity and mass. $f_{e}$ is electron velocity distribution function, which is assumed to be Maxwellian.
@@ II.1 The morphology of debris-induced nonlinearities
Before we investigate the nonlinear waves and structures induced by a charged debris in a PINI plasma, it is of relevance to discuss a bit about the fundamental physics of these nonlinearities in an $e$-$i$ plasma. We now know that the response of an $e$-$i$ plasma to an embedded charged debris very well depends on the nature of the charge of the debris [22]. It has been only recently shown that the formation of bright pinned solitons in the ion-acoustic regime due to the presence of negatively charged debris is basically a manifestation of the trapped ions in phase-space vortices formed due to ion-ion counter streaming instability (IICSI) [19]. Besides, proper spatial resolution of the pinned solitons requires a certain relative velocity between the debris and the plasma, though higher relative velocities may cause the pinned solitons to disappear [17, 19]. In contrast, a positively charged debris creates an ion hole, exciting an ion-acoustic wave (IAW), which propagates away from the site of the debris. However, when there is a relative motion between the debris and the plasma, the IAW becomes a dispersive shock wave (DSW) in the precursor region [22]. We should, however be careful, while referring to these so-called ‘pinned’ solitons. Though they are referred to as ‘solitons’, they are characteristically different from true solitons as a true soliton does not evolve in time, retains its shape, and is asymptotically stable, which is produced by an integrable system. On the other hand, pinned solitons, produced by external charged debris are not true solitons as they are not asymptotically stable and produced by non-integrable systems.
@@ II.2 Mathematical reduction
We note that the above plasma model can be analyzed both mathematically and numerically. Typically, in this class of problems, one uses a scale transformation and reduces it to a pseudo potential (or Sagdeev potential) form [26] through the Poisson equation, which guarantees a soliton solution, though not necessarily a ‘${\rm sech}$’-type solution as one would have obtained through a KdV equation [27]. The other two approaches would be a reductive perturbation analysis to reduce the system to a KdV equation or a nonlinear Schrödinger equation (NLSE) [28]. However, with an external charged debris term, the model is simply not amenable to pseudo-potential analysis and that leaves us with other two options.
Employing the reductive perturbation method with the external charged debris term, we end up with either a forced-KdV (fKdV) or a forced-NLSE equation. At this point, we should emphasize that numerically we are going to a use flux-corrected transport (FCT) method (as mentioned before) to simulate the dynamics, which is very efficient in detecting sharp discontinuities that can result from the debris-induced perturbation. Also the FCT formalism does not employ any approximations and solves the full set of equations, which in principle should provide us with both KdV-like as well as NLSE-like modulated wave solutions. However, the single hump perturbation that we expect from an external charged debris, the numerical solution is most likely to favor KdV-like solution which is also closer to the real-world and observable dynamics. In what follows, we shall therefore analyze the above model mathematically, using the fKdV equation only.
@@ III Forced-KdV dynamics
where $\varepsilon$ is the expansion parameter, which should be $\ll 1$. The expansion parameter $\varepsilon$ is not a physical parameter of the system in the sense that it does not come from measurement of any physical quantity and it needs to be interpreted in the context of the numerical solution of the resultant fKdV equation (see Appendix A for a possible definition). One question related to the order of perturbation for the external charged debris may naturally arise here. In this context, two points to be noted here – that we usually use the KdV-type equation for weak nonlinearity and how strong a perturbation an external charged debris may induce in the bulk plasma. Our assumption of using a second order perturbation for the debris is quite in line with the weak nonlinearity assumption on the basis of which the fKdV equation is derived. As far as the strength of perturbation is concerned, we shall see in subsequent sections that the experimentally observed nonlinearities are quite close to our numerical solutions (obtained via FCT simulation) as well as the fKdV solutions, which demonstrates that our assumption of introducing the external charged debris as second order perturbation is indeed consistent (an explanation justifying this is given in Appendix A). We should further note that the reductive perturbation theory essentially requires that all external perturbations must not exist at the first order or else the theory would fail (see Appendix B for a simple explanation). Besides, using a second order perturbation for debris leaves us with the simplest form of the fKdV equation.
@@
As mentioned before, we now provide an explanation about our ‘dissipation-less’ assumption of the plasma model. The dissipation in this case may come in two flavors – through collisions and Landau damping. There are examples of work involving multicomponent plasma (not necessarily PINI plasma) with collisions as well as with Landau damping term [29]. While the inclusion of different collisional terms in the fluid model is quite straightforward [30, 31], the same is not true in the case of Landau damping. However, as shown by VanDam [32], it is indeed possible to include a dissipation-like term arising out of wave-particle resonance (Landau damping) in a KdV-like formalism. In Section III-A.1, while comparing our fKdV solutions with ion-beam-induced nonlinearities, we shall see that the $e$-$i$ collision time $\tau_{ei}\sim 3.5\,{\rm milli}\,{\rm second}$. However, the ion-acoustic time scale, which is the characteristic time scale of this work, is $\tau_{{\rm IA}}\sim 0.03\,{\rm milli}\,{\rm second}$. So, we can probably safely neglect collisional dissipation. Coming back to now Landau damping, VanDam [32] has shown that a dissipation-like term can be incorporated into the KdV equation with the linear Landau damping rate $\gamma=\tilde{\gamma}|k|$ (see Appendix C) where
@@ III.1 Positively charged external debris
Figure 1: Nonlinear wave for a flowing PINI plasma with a positively charged external debris. Panels (a)-(f) correspond to debris velocities $v_{d}=0.5,1.0,1.1,1.5,1.8$, and $2.1$, respectively. All panels are drawn in the rest frame of the debris, which is in the middle of a panel, denoted by a vertical dashed line. The shaded vertical regions indicate the extent of the charged debris. The shaded envelopes on the right hand panels indicated the envelopes of the corresponding dark pinned solitons.
determined by $\Delta$ [16]. We note that the external debris is introduced as a perturbation with $\rho_{0}$ denoting the strength. In all the cases (unless indicated otherwise), we have used a perturbation with $\rho_{0}\sim 0.5$ which is equivalent to a perturbation $\equiv 50\%$ of the equilibrium density. Regarding the perturbation amplitude, we would like to note that the magnitude of perturbation of $\sim 50\%$ of the equilibrium density may not look like a weak perturbation. However, when we translate it to the nonlinearity regime, the perturbation indeed has introduced a nonlinearity, which the
Figure 2: Nonlinear wave for a flowing plasma with a negatively charged external debris. Panels (a)-(f) correspond to debris velocities $v_{d}=0.5,1.0,1.1,1.5,1.8$, and $2.1$ respectively. All other particulars are same as in Fig.1. The shaded vertical regions indicate the extent of the negatively charged debris.
fKdV equation can handle (see Appendix A). This fact is verified through FCT simulation (see Section IV), where we have estimated the expansion parameter $\varepsilon$ from simulation to be about $\sim 0.7$ with simulation results closely agreeing our fKdV results. In the figure, the shaded region shows the extent of the debris charge distribution centered in the middle of the frame indicated by the dashed vertical line. The debris move with a velocity $v_{d}$ from right to left.
In Fig.1(a), we observe that for low (subsonic) $v_{d}$, the debris excites bright ion-acoustic solitary wave which move away from the debris in the precursor region. We also observe that the velocities of these solitary waves increase from being subsonic to becoming supersonic as $\delta$ increases from $1.02$ to $1.8$. The changing shape of the soliton-like structure for increasing $\delta$ can be attributed to the increasing negative ion density (with higher $\delta$) which tend to accelerate from left to right as a response to the positive perturbation at the debris site. We further note that in the production of a PINI plasma, experimentalists usually use an electronegative gas in an $e$-$i$ plasma to create negative ions at the expense of less warm electrons. So, an increase of $\delta$ causes an increase in negative ion density in a PINI plasma and the increase of velocity of the soliton can be attributed to the increase in negative ion density. However, as debris velocity increases and becomes sonic, higher $\delta$ perturbations slow down in comparison to lower $\delta$ perturbations which ultimately forms dark pinned solitons as debris velocity increases more, as shown in the three panels on the right in the figure, Fig.1(d), (e), and (f). As $v_{d}$ increases, pinned solitons tend to form at the debris position for higher $\delta$ whereas low $\delta$ perturbations never get to form pinned solitons and remain predominantly a positive nonlinearity in the form of a DSW in the precursor region.
The assumption of Gaussian-shaped distribution for the charge density of a debris of size $\gg\lambda_{D}$ can be fairly ascertained, if the parameter $\beta_{e}\ll 1$ (for negatively charged debris), with
@@
where $e$ is the electronic charge and $R$ is the radius of a spherical debris [33]. This result is rigorously proved for impurity particles in a plasma using the orbit motion limited (OML) theory [34] and numerical simulation, which can be readily interpreted for charged debris as well. This can also be extended to debris with positive charge. If we now consider the case of charged debris in space plasma in Low Earth Orbit (LEO), the size of the debris can vary from being very small to very large, but anything larger than a few Debye length is capable of inducing nonlinear structures. Considering typical parameters for LEO plasma [35, 36], the value for $\beta_{e}$ comes out to be $\sim 10^{-6}$ even for $R\sim\lambda_{D}$. For ions (positively charged debris), this comes out to be $\beta_{i}\sim 10^{-3}$. In laboratory experiments involving debris $\beta_{e}\sim 10^{-5}$ assuming debris of size $\sim\lambda_{D}$ [20].
These results are entirely in conformity with what we know about formation of debris-induced pinned solitons and DSWs. We now know that in an $e$-$i$ plasma the pinned solitons formed due to moving negatively charged debris is basically a manifestation of the ions trapped in the phase space vortices formed by the IICSI [19], whereas a positively charged debris induces a DSW in the precursor region. So, for $\delta\sim 1$ (very less concentration of negative ions), the plasma behaves like the usual $e$-$i$ plasma and no pinned solitons can form for a moving positively charged debris. However for $\delta>1$, the negative ions are attracted by the positively charged debris causing a IICSI-like phenomena for negative ions, leading to the formation of phase space vortices which in turn cause formation of dark pinned solitons at the debris position.
@@ III.1.1 Comparison with ion-beam-induced nonlinearities
We now discuss the first case where under certain circumstances, a positive ion-beam in a PINI plasma mimics the responses of a debris-induced nonlinearities. In this experiment, the positive ion-beam is produced in situ by accelerating a part of the ion population in the plasma so that overall quasi-neutrality is maintained throughout [11]. This case is also interesting as there are several investigations being carried out by different authors to analyze the formation of nonlinear structures in such plasmas [12, 13]. Let us consider such a three-component PINI plasma, consisting of positive ion $(n_{+})$, negative ion $(n_{-})$, and electron $(n_{e})$, in the presence of a positive ion beam $(b^{+})$, as considered in the experimental work by Nakamura [11] (N99 hereafter). Note that the whole system is to be considered quasi-neutral,
@@
where $\ln\Lambda$ is the Coulomb logarithm $\sim 10-12$. For $n_{e}\sim 10^{7}\,{\rm cm}^{-3}$ [11], we have $\tau_{ei}\sim 3.5$ milli second. We now argue that as $\tau_{ei}\gg\tau_{{\rm IA}}$, it can be safely assumed that in the ion-acoustic timescale, the plasma will not be able to thermalize. As a result, a positive ion beam propagating through the plasma will not be able to thermalize with the bulk plasma and for all practical purposes, the beam will behave as an external debris. It has been observed in N99 that below a certain critical value for beam velocity, increasing $r$ (higher $\delta$) causes the nonlinearity to become negative from positive. However, beyond the critical value of the beam velocity, the nonlinearity remains positive irrespective of the initial pulse (which is reported only at higher $\delta$ in N99). In our case with debris as well, we see a similar behavior as evident from the panels in the left column of the figure i.e. Fig.1(a), (b), and (c). We can see that when debris velocity is $\apprle 1.1$, for low $\delta$ $(\sim 1)$, there is a steepening in the leading edge indicating positive nonlinearity while for higher $\delta$, the steepening is on the falling edge, which indicates negative nonlinearity. We note that the ‘positive nonlinearity’ in the case of a compressive or bright soliton-like structure is referred with respect to the nonlinear term $A\phi^{(1)}\partial_{\xi}\phi^{(1)}$ [see Eq.(16)] in a usual KdV equation (without the debris term), which is responsible for steepening of the leading edge of a soliton-like structure. On the other hand, when the steepening occurs in the falling edge, it is usually termed as ‘negative nonlinearity’. When debris velocity increases [Fig.1(f)], the negative nonlinearity which was observed for say $\delta=1.5$ and with low $v_{d}$, decreases and gets balanced by dispersion giving rise to a bright KdV soliton-like pulse. These results are also confirmed through our FCT simulation, which are being presented in Section IV.
@@
Figure 3: FCT simulation results of propagation of nonlinear wave in a PINI plasma with an embedded external positively charged debris. These results are equivalent to the theoretical cases obtained through fKdV solution shown in Fig.1. The left panel (a) in the above figure corresponds to Fig.1(c) and the right panel (b) above corresponds to Fig.1(d). These results are plotted at the end of $\tau=100\equiv\varepsilon^{3/2}t$.